4. The cost of preferred stock  

Firms that carry preferred stock in their capital mix want to not only distribute dividends to the company's common n stockholders but also maintain credibility in the capital markets so that they can raise additional funds in the future and avoid potential corporate raids from preferred stockholders.

Consider the case of Purple Lemon Shipbuilders Inc.

The CFO of Purple Lemon Shipbuilders Inc. has decided that the company needs to raise additional capital. It can sell preferred stock paying an annual $6 dividend per share for $100 per share; however, it will incur a flotation cost of 1.3% per share.

1) After it pays the underwriter, Purple Lemon Shipbuilders Inc. will receive__________ from each share of preferred stock that it issues.

2) Based on this information, Purple Lemon Shipbuilders Inc.'s cost of preferred stock is __________

3) When raising funds by issuing new preferred stock, the company will incur an underwriting, or flotation, cost that __________the cost of preferred stock.

4) Because the flotation cost is usually expressed as a percentage of prices of each share, the difference between the cost of preferred stock with and without flotation cost is__________ enough to not ignore.
Answer #1

1)

floatation cost , f = 1.3% per share = 0.013

f = 0.013*p = 0.013*100 = 1.3

dividend per preferred share = $6

price of preferred stock , p = $100

if this is the price including cost of prferred stock,

then the company actually receives an amount per share P= p*(1-f) = 100*(1-0.013) = $98.7

2)

Based on this information, Purple Lemon Shipbuilders Inc.'s cost of preferred stock is : Dividend per share/P = 6/98.7 = 0.0607902 = 6.07902% or 6.08% ( after rounding off to 2 decimal places)

3)

When raising funds by issuing new preferred stock, the company will incur an underwriting, or flotation, cost that Increases the cost of preferred stock

4)

Because the flotation cost is usually expressed as a percentage of prices of each share, the difference between the cost of preferred stock with and without flotation cost is  Big ( or high) enough to not ignore